OpenStudy (anonymous):

If u = <-5, -3> and v = <-4, 6>, find u · v. 38 2 -2 -38

OpenStudy (amorfide):

u.v=(u1v1 + u2v2)

OpenStudy (michele_laino):

by definition of scalar product, we have: \[u \cdot v = \left( { - 5} \right) \times \left( { - 4} \right) + \left( { - 3} \right) \times 6 = ...?\]

OpenStudy (anonymous):

20x

OpenStudy (anonymous):

+

OpenStudy (michele_laino):

\[u \cdot v = \left( { - 5} \right) \times \left( { - 4} \right) + \left( { - 3} \right) \times 6 = 20 - 18 = ...?\]

OpenStudy (anonymous):

2

OpenStudy (michele_laino):

that's right!